OpenStudy (lynn.caroline):

If f(x)=7x-5x^2, then find each of the following: A.) f(3k-7) B.) f(x+h) C.) f(-4k^5)

TheSmartOne (thesmartone):

You simply plug in 3k-7 in place of every x and simplify f(x) = 7x - 5x^2 \(f(\color{red}{3k - 7}) = 7(\color{red}{3k-7}) - 5(\color{red}{3k-7})^2\)

OpenStudy (lynn.caroline):

Okay that makes sense! Will you check my answers when I'm done?

OpenStudy (lynn.caroline):

A I got -14

TheSmartOne (thesmartone):

Sure! :)

TheSmartOne (thesmartone):

We have variables, how did you get rid of them? Can you show your work so I can see where you went wrong? (:

OpenStudy (lynn.caroline):

I used my calculator, guess that doesn't work lol. Let me work it out

OpenStudy (lynn.caroline):

Okay I'm stumped

TheSmartOne (thesmartone):

to expand (3k-7)^2 use the formula \(\Large (a-b)^2 = a^2 - 2ab + b^2\) and to open up 7(3k-7) use the distributive property which states a(b-c) = ab - ac

TheSmartOne (thesmartone):

and after you expand (3k-7)^2 don't forget to distribute the -5 ;o

OpenStudy (lynn.caroline):

I ended up with f(3k-7)= 21k + 76. I feel like that's not right

OpenStudy (lynn.caroline):

Do I need to use that formula for 7(2k-7)?

TheSmartOne (thesmartone):

no, you need to use the distributive property for 7(3k-7) psst, it's 3k not 2k and that's not right either can you show your work? :p

OpenStudy (lynn.caroline):

I think I'm doing this all wrong :( Okay f(3k-7) = 7(3k-7) -5(3k-7)^2 (3+7)^2 = 3 - 2(3)(-7)-7^2 10^2 = 3 + 42 - 49 100/-4 = -4/-4 -25 -5(-25) = 125 Then 7(3k-7) 21k - 49 -49+125+21k 76 + 21k

TheSmartOne (thesmartone):

"(3+7)^2 = 3 - 2(3)(-7)-7^2" it's 3k -7, and I don't know what formula you're using o.o try this (3k-7)^2 = (3k)^2 - 2(3k)(7) + 7^2 what do you get?

OpenStudy (lynn.caroline):

3k^2 - 6x + 35

TheSmartOne (thesmartone):

\((ab)^m = a^mb^m\) so what is \( (3k)^2 = ?\) and you changed k to x and you didn't multiply 2 * 3 and 7 properly and 7^2 is not 35

OpenStudy (lynn.caroline):

I thought 3k^2 just stayed like that? My bad in multiplied 7 by -2 :( Here's what I did (3k)^2 - 2(3k)(7) + 7^2 3k^2-6k -14 +49 3k^2 - 6k + 35

TheSmartOne (thesmartone):

7^2 = 7 * 7 = 49 \( (3k)^2 = 3^2k^2 = 9k^2\) and 2(3k)(7) = 42k sooooo (3k-7)^2 = 9k^2 - 42k + 49 do you understand? o:

OpenStudy (lynn.caroline):

That makes a lot more sense. So sorry for the confusion, math isn't my strong subject /:

TheSmartOne (thesmartone):

so what is -5(9k^2 - 42k + 49) distribute the -5 an extension of the distributive property a(b - c + d) = ab - ac + ad

OpenStudy (lynn.caroline):

-45k^2 + 210 -245?

TheSmartOne (thesmartone):

yesss, but you dropped the k after 210 and what is 7(3k-7) = ?

OpenStudy (lynn.caroline):

Oops typo! 7(3k-7) 21k - 49?

TheSmartOne (thesmartone):

yes, now 21k - 49 -45k^2 + 210 -245 = ?? simplify it by adding/subtracting like terms

OpenStudy (lynn.caroline):

-45k^2 + 21k -84

TheSmartOne (thesmartone):

oops, my bad 21k - 49 -45k^2 + 210k -245 = ?? I forgot to put the k after 210 ;p

OpenStudy (lynn.caroline):

You're good! So -45k^2 + 231k - 294

TheSmartOne (thesmartone):

bingo!

OpenStudy (lynn.caroline):

You're amazing! Would I set b up the same way

TheSmartOne (thesmartone):

yes, instead of 3k-7, you would use x + h

OpenStudy (lynn.caroline):

So f(x+h) = 7(x+h) - 5(x+8)^2 Start with (x+h)^2 = x^2 - 2(x)(h) + h^2 right?

TheSmartOne (thesmartone):

yep and then distribute the -5

OpenStudy (lynn.caroline):

Okay I got -5x^2 - 5h^2 + 10xh

TheSmartOne (thesmartone):

so we have so far 7(x+h) - 5(x+8)^2 7(x + h) -5x^2 - 5h^2 + 10xh now do 7(x + h)

OpenStudy (lynn.caroline):

7x + 7h

TheSmartOne (thesmartone):

7x + 7h -5x^2 - 5h^2 + 10xh now simplify (:

OpenStudy (lynn.caroline):

Can that be simplified?

OpenStudy (lynn.caroline):

The only thing I could think to do would be 24xh -5h^2 - 5x^2

TheSmartOne (thesmartone):

it can't be simplified :b you can't add h + h to make h^2 only h * h = h^2 h + h = 2h but we can't simplify anymore :)

OpenStudy (lynn.caroline):

Okay so it stays how it was?

TheSmartOne (thesmartone):

yeah xD

OpenStudy (lynn.caroline):

You tricked me lol. So for c do we still use the whole (a-b)^2=a^2-2ab+b^2 thing?

TheSmartOne (thesmartone):

nope, C is much more simpler xD

OpenStudy (lynn.caroline):

Okay so 7(-4k^5) = -28k^5. Do the same to the other side?

TheSmartOne (thesmartone):

yeah, but we first have to square it :)

OpenStudy (lynn.caroline):

400k^10?

TheSmartOne (thesmartone):

nooo \((-4k^5)^2 = (-4)^2(k^5)^2\) hint: \( (a^b)^c = a^{bc}\)

OpenStudy (lynn.caroline):

16 and k^10? Or would it be k^7?

TheSmartOne (thesmartone):

k^10 is correct and then 5 * 16k^10 = ?

OpenStudy (lynn.caroline):

80k^10!

OpenStudy (lynn.caroline):

So the answer is 480k^10?

TheSmartOne (thesmartone):

we can't add -28k^5 and 80k^10 and not sure how you got 480 o;

OpenStudy (lynn.caroline):

Shoot idk where I got 400k^2 from lol. I added that. So 80k^10-28k^5

TheSmartOne (thesmartone):

xD

OpenStudy (lynn.caroline):

Thank you so so much for explaining that! Sorry it took so long

TheSmartOne (thesmartone):

It's alright! Have a great day (: